Let f(x) = 3x – 6 and g(x) = 6x + 5. Write a formula for each of the following functions and then simplifye. fg= Preview f. ff= Preview
Answer #1

f 0 = 34-6, g(= 6xts fg = (가) (6u+S) 182 15M -361 - 30 18x² 2 -30 (1) 과 다. - 9 36-2(31)(6) 무 무 36-367